OpenStudy (oswaldmurphy):

Part A: Maria rented a coat at $285 for 3 days. If she rents the same coat for 6 days, she has to pay a total rent of $510. Write an equation in the standard form to represent the total rent (y) that Maria has to pay for renting the coat for x days. (4 points) Part B: Write the equation obtained in Part A using function notation.(2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

OpenStudy (oswaldmurphy):

@TheSmartOne

OpenStudy (oswaldmurphy):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

x = number of days y = total cost `Maria rented a coat at $285 for 3 days.` one ordered pair is (3,285) `she rents the same coat for 6 days, she has to pay a total rent of $510` another ordered pair is (6,510)

jimthompson5910 (jim_thompson5910):

find the equation of the line through (3,285) and (6,510)

OpenStudy (oswaldmurphy):

510 - 285 / 6 - 3 ?

jimthompson5910 (jim_thompson5910):

use parenthesis to properly group things up (510-285)/(6-3)

jimthompson5910 (jim_thompson5910):

but yes

jimthompson5910 (jim_thompson5910):

that will give you the slope

OpenStudy (oswaldmurphy):

Its asking for standard form, ax + by = c, how would I do that?

jimthompson5910 (jim_thompson5910):

what slope did you get?

OpenStudy (oswaldmurphy):

75

jimthompson5910 (jim_thompson5910):

so pick any point, say (3,285) and use this point along with the slope to plug them into point-slope form given point: (x1,y1) = (3,285) x1 = 3 y1 = 285 slope: m = 75 \[\Large y - y_1 = m(x - x_1)\] \[\Large y - 285 = 75(x - 3)\] \[\Large y - 285 = 75x - 225\] get this into the form Ax+By = C

OpenStudy (oswaldmurphy):

How?

jimthompson5910 (jim_thompson5910):

get the x and y terms to one side. Get everything else to the other side

OpenStudy (oswaldmurphy):

So -60 = 75x - y?

jimthompson5910 (jim_thompson5910):

or 75x - y = -60

jimthompson5910 (jim_thompson5910):

it's in the form Ax+By = C where A = 75 B = -1 C = -60

OpenStudy (oswaldmurphy):

Ok, so that is the answer to a?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

I recommend you plug in each point, (3,285) and (6,510), to confirm that this is actually the line that goes through both points

jimthompson5910 (jim_thompson5910):

I'll show you how to check (3,285) x = 3 y = 285 75x - y = -60 75*3 - y = -60 ... plug in x = 3 75*3 - 285 = -60 ... plug in y = 285 225 - 285 = -60 -60 = -60 ... works

OpenStudy (oswaldmurphy):

I could also use Desmos correct?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

It's always a good idea to practice plugging in the points just in case you're not able to use a graphing device on an exam

OpenStudy (oswaldmurphy):

Vary steep XD.

OpenStudy (oswaldmurphy):

jim its an online course so I should be fine, but if not I will check. Thanks! How do I do B?

jimthompson5910 (jim_thompson5910):

you'll have to adjust the window

jimthompson5910 (jim_thompson5910):

Part B) solve `75x - y = -60` for y. Then replace y with f(x)

OpenStudy (oswaldmurphy):

-75x from both sides. -y = -60 - 75x. Divide it all by -1. y = 60 + 75x f(x) = 60 + 75x

jimthompson5910 (jim_thompson5910):

yes or f(x) = 75x+60

OpenStudy (oswaldmurphy):

Oh ok, Ill start putting it into correct format.

OpenStudy (oswaldmurphy):

Now C. Would It simply be (3, 285) and (6, 510)?

jimthompson5910 (jim_thompson5910):

Plot those two points on the same coordinate axis. Then draw a straight line through the points. Make sure the line extends as far as you can in both directions.

OpenStudy (oswaldmurphy):

I am not supposed to do that, it says describe how to.

jimthompson5910 (jim_thompson5910):

I just described it

OpenStudy (oswaldmurphy):

Oh ok I thought you were telling me to do that XD

jimthompson5910 (jim_thompson5910):

You can if you want, but the two ideas are very similar

OpenStudy (oswaldmurphy):

A: 75x - y = -60 B: f(x) = 75x + 60 C: Plot those two points on the same coordinate axis. Then draw a straight line through the points. Make sure the line extends as far as you can in both directions.

OpenStudy (oswaldmurphy):

Look good?

jimthompson5910 (jim_thompson5910):

be sure to elaborate on what you mean by "those two points" for part C

OpenStudy (oswaldmurphy):

Plot those two points (3, 285) and (6, 510) on the same coordinate axis. Then draw a straight line through the points. Make sure the line extends as far as you can in both directions.

jimthompson5910 (jim_thompson5910):

good

OpenStudy (oswaldmurphy):

Thanks I gave you a medal and a fan and a testi